### v1

# DART: Difficulty-Adaptive Reasoning Truncation for Efficient Large Language Models

Abstract
[…]
1 Introduction
The emergence of chain-of-thought (CoT) reasoning has marked a significant advance in enhancing the problem-solving abilities of LLMs by decomposing complex questions into intermediate steps (Wei et al., 2022; Kojima et al., 2022). Despite its effectiveness, the conventional CoT paradigm exhibits a critical inefficiency: it typically generates reasoning chains of a fixed, often excessive, length regardless of the inherent difficulty of the problem at hand (Chen et al., 2024; Fan et al., 2025; Sui et al., 2025). This “one-size-fits-all” approach results in substantial computational redundancy, increasing inference latency and resource consumption—a major bottleneck for deploying LLMs in applications.
Figure 1: Overall workflow of the proposed DART framework.
Adaptive reasoning, which aligns computational effort with problem difficulty, offers a promising path toward efficiency. Recent attempts to achieve such adaptability have largely relied on reinforcement learning (RL) frameworks, training models to penalize unnecessary reasoning length while preserving accuracy (Arora & Zanette, 2025; Ling et al., 2025; Zhang et al., 2025b; Shen et al., 2025). However, these RL-based methods remain unstable and heavily reward-dependent, suffering from high training difficulty and limited generalizability. Alternative approaches such as knowledge distillation focus on generating shorter reasoning chains (Yu et al., 2024; Kang et al., 2025; Xia et al., 2025), yet they produce statically compressed CoTs that lack dynamic adaptability. The key challenge is therefore to enable difficulty-aware reasoning in a stable and general manner, without relying on complex RL pipelines.
[…]
2 Related Work
Reasoning Chain Compression and Distillation. To address the computational overhead of lengthy reasoning chains, several compression and distillation techniques have been proposed. Knowledge distillation methods train smaller student models to mimic the reasoning processes of larger teacher models (Yu et al., 2024; Kang et al., 2025; Xia et al., 2025). These approaches typically produce statistically compressed reasoning chains that maintain a fixed length across all problems. Prompt-based compression techniques (Xu et al., 2025; Han et al., 2024) attempt to generate shorter reasoning chains through specialized prompting strategies, but often struggle with accuracy preservation on complex problems. Our work differs by introducing dynamic compression that adapts reasoning length to problem difficulty, achieving better efficiency-accuracy trade-offs than static compression methods.
Adaptive Reasoning Methods. Adaptive reasoning aims to dynamically adjust the reasoning process based on problem characteristics. Reinforcement learning approaches (Arora & Zanette, 2025; Yu et al., 2025; Shen et al., 2025; Zhang et al., 2025a) train length policies to decide when to halt reasoning, but face challenges with training stability and reward engineering. Prompt-based adaptive methods (Han et al., 2024) use heuristic rules to control reasoning length, but lack learning-based optimization. Unlike these approaches, DART employs a novel supervised learning framework that learns optimal reasoning length policies from automatically curated data, avoiding the instability of RL methods while maintaining architectural flexibility.
3 Method
3.1 Motivation
The core motivation behind our work stems from a fundamental observation on the relationship between the amount of reasoning (measured in the number of tokens generated) and the final reasoning accuracy.
To quantify this, we conducted a preliminary experiment on a subset of 10 problems from the MATH-500 (Lightman et al., 2023) dataset. Using our model fusion technique (detailed in Section 3.3), we generated reasoning chains for these problems across a spectrum of lengths, controlled by the fusion coefficient $\alpha$. For each problem, we evaluated the correctness of the answer yielded by chains of varying lengths. We then analyzed the aggregate accuracy as a function of the average number of tokens in the reasoning chains.
Figure 2: The trade-off between reasoning chain length and accuracy, illustrated with 10 problems from the MATH-500 dataset.
The key observation is that the resulting accuracy-vs-tokens curve exhibits a pronounced Sigmoid (S-shaped) pattern, as illustrated in Fig. 2. This curve reveals three distinct phases: i) an initial low-accuracy plateau with insufficient reasoning steps, ii) a rapid ascent where token increases significantly boost accuracy (the “sweet spot”), and iii) a final plateau of diminishing returns where additional tokens yield minimal gains.
This Sigmoid relationship highlights a critical inefficiency in standard CoT generation: for given problem, there exists an optimal reasoning length that is sufficient to achieve the highest accuracy, and generating anything beyond this point is wasteful.
However, this optimal length is not universal; it is tied to the difficulty of the individual problem. Simple problems may reside on the upper plateau, requiring only a few steps, while complex ones may need to be in the rapid ascent phase to be solved correctly. Current methods—whether generating long chains or distilling fixed short chains—operate at a single point on this curve. They force all problems to be processed with a one-size-fits-all reasoning budget, inevitably leading to inefficiency (for easy problems) or inadequacy (for hard ones).
Therefore, the goal is not to find a single best average length, but to enable a model to dynamically locate the minimal sufficient point on this curve for each input problem. This is the definition of adaptive thinking that our method, DART, is designed to achieve.
3.2 Step 1: Distilling Short CoTs
This initial step aims to establish the two core model components for the DART framework: a compact Base Model that possesses inherent long-chain reasoning capability, and a Distilled Model derived from it that produces concise reasoning chains. This distillation process is crucial for creating the “short-chain” endpoint of the reasoning spectrum, which will be interpolated with the powerful pre-existing “long-chain” base model.
Base Model ($M_{\mathrm{base}}$). We select an efficient, open-source base model that natively supports chain-of-thought reasoning, such as DeepSeek-R1 (Guo et al., 2025) series or Qwen3 (Yang et al., 2025) series. This model, denoted as $M_{\mathrm{base}}$, is employed as-is for its ability to generate detailed, step-by-step reasoning. It serves as the efficient backbone whose behavior we aim to adapt, representing the long-chain reasoning style.
Short CoT Data Distillation ($\mathcal{D}_{\mathrm{short}}$).
We construct the short-chain dataset by compressing existing high-quality CoT data. Using a powerful distillation teacher model (e.g., DeepSeek-R1), we shorten each long reasoning chain $\mathrm{CoT}_{i}^{\mathrm{long}}$ to a concise version $\mathrm{CoT}_{i}^{\mathrm{short}}$ while preserving logical correctness:
$\mathrm{CoT}_{i}^{\mathrm{short}}=M_{\mathrm{distillation-teacher}}(\mathrm{Prompt}_{\mathrm{compress}}(x_{i},y_{i},\mathrm{CoT}_{i}^{\mathrm{long}}))$ (1)
The resulting dataset retains the original question $x_{i}$ and answers $y_{i}$ but contains significantly shortened reasoning paths.
Training the Distilled Model ($M_{\mathrm{distilled}}$). The Distilled Model ($M_{\mathrm{distilled}}$) is created by performing supervised fine-tuning (SFT) on the original $M_{\mathrm{base}}$ using the compressed dataset $\mathcal{D}_{\mathrm{short}}$. This model learns to produce accurate answers with minimal reasoning, representing the “short-chain” endpoint. The learning objective is to adapt the base model’s reasoning behavior to generate concise chains:
$\mathcal{L}_{\mathrm{distilled}}=-\sum\nolimits_{t}\log P(w_{t}^{\mathrm{short}}|x,w_{<t}^{\mathrm{short}})$ (2)
By the end of this step, we have two models that share an architectural origin but possess distinct “reasoning styles”. $M_{\mathrm{base}}$ embodies thoroughness at the cost of efficiency, while $M_{\mathrm{distilled}}$ embodies extreme efficiency. The interplay between these two models enables the adaptive capabilities developed in the subsequent steps.
3.3 Step 2: Creating a Model Spectrum via Fusion
[…]
Fusion Process. For any given $\alpha$, the parameters $\theta_{\alpha}$ of the fused model $M_{\mathrm{\alpha}}$ are calculated as a weighted linear combination of the parameters of the base and distilled models:
$\theta_{\alpha}=(1-\alpha)\cdot\theta_{\mathrm{base}}+\alpha\cdot\theta_{\mathrm{distilled}}$ (3)
where $\theta_{\mathrm{base}}$ are the parameters of $M_{\mathrm{base}}$, $\theta_{\mathrm{distilled}}$ are the parameters of $M_{\mathrm{distilled}}$, and $\alpha$ is the fusion coefficient. The resulting model $M_{\mathrm{\alpha}}$ is instantiated with the parameters $\theta_{\alpha}$.
Interpretation of the Fusion Spectrum. The coefficient $\alpha$ dictates the ”reasoning identity” of the fused model: i) When $\alpha=0$, $M_{\mathrm{\alpha=0}}=M_{\mathrm{base}}$. This model generates the longest, most detailed chains. ii) When $\alpha=1$, $M_{\mathrm{\alpha=1}}=M_{\mathrm{distilled}}$. This model generates the shortest, most compressed chains. iii) When $0<\alpha<1$, the model $M_{\alpha}$ blends the behaviors of its parents. As $\alpha$ increases from 0 to 1, the generated reasoning chains become progressively shorter and more concise, effectively traversing the accuracy-vs-tokens curve introduced in Section 3.1.
3.4 Step 3: Curating the Adaptive Training Data
The model spectrum $M_{\alpha}$} generated in the previous step provides a diverse set of reasoning strategies for any given problem. The goal of this step is to automate the construction of a high-quality training dataset $\mathcal{D}_{\mathrm{adaptive}}$ where each problem is paired with its optimal reasoning chain—defined as the shortest chain that leads to a correct answer. This dataset will directly teach the final model the skill of adaptive thinking.
Data Curation Pipeline. For each problem $x_{i}$ in our training set, we execute the following pipeline:
1. Reasoning Generation: We input $x_{i}$ into every model $M_{\alpha}$ in our sampled spectrum. Each model generates a reasoning chain $\mathrm{CoT}_{i}^{\mathrm{\alpha}}$ and a predicted answer $y_{i}^{\alpha}$.
2. Answer Verification: We verify the correctness of each predicted answer $y_{i}^{\alpha}$ against the ground-truth answer $y_{i}$ using a task-specific criterion (e.g., exact match for mathematical answers, predefined evaluation metrics for other reasoning tasks).
3. Optimal Chain Selection: From all models that produced the correct answer ($y_{i}^{\alpha}=y_{i}$), we select the reasoning chain $\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}}$ from the model with the largest value of $\alpha$ (i.e., the model biased towards the shortest chains). Formally: $\alpha^{*}=\min\{\alpha\in S\mid y_{i}^{\alpha}=y_{i}\}$, where $S$ is the set of sampled $\alpha$ values. The corresponding $\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}}$ is deemed the optimal adaptive chain for problem $x_{i}$.
4. Data Assignment: The tuple $(x_{i},y_{i},\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}})$ is added to the new adaptive training dataset $\mathcal{D}_{\mathrm{adaptive}}$.
Handling Edge Cases. If no model in the spectrum produces the correct answer, the problem $x_{i}$ is excluded from $\mathcal{D}_{\mathrm{adaptive}}$. This ensures the quality of the training data. If multiple models with the same $\alpha$ value produce the correct answer, the shortest generated chain among them is selected, providing a further refinement.
Theoretical Justification and Outcome. This selection protocol is a data-driven implementation of the motivation described in Section 3.1. For each problem, it identifies the operational point near the ”elbow” of the sigmoid curve—the point where accuracy is achieved with minimal computational cost. The resulting dataset $\mathcal{D}_{\mathrm{adaptive}}$ no longer contains chains of a fixed length. Instead, it is a collection of difficulty-labeled examples by proxy; the length of the chain $\mathrm{CoT}_{i}^{\mathrm{\alpha^{*}}}$ represents the complexity of the problem $x_{i}$.
By learning from these optimal (problem, chain) pairs, a model can be trained to emulate this optimal selection process dynamically at inference time. The final dataset $\mathcal{D}_{\mathrm{adaptive}}=\{(x_{i},y_{i},\mathrm{CoT}_{i}^{\mathrm{opt}})\}_{i=1}^{M}$ is the key resource for training our adaptive model in the final step.
3.5 Step 4: Training the Adaptive Model
The final and crucial step of the DART framework is to distill the collective knowledge of the model spectrum and the optimal adaptive policy embodied in $\mathcal{D}_{\mathrm{adaptive}}$ into a single, efficient, and standalone model, denoted as $M_{\mathrm{adaptive}}$. This model is designed to intrinsically learn the mapping from problem difficulty to reasoning length, enabling it to generate the minimally sufficient chain-of-thought autonomously during inference, without relying on the cumbersome process of generating multiple chains from a spectrum of models.
Training Objective. We initialize $M_{\mathrm{adaptive}}$ from the same pre-trained checkpoint as the base and distilled models. The model is then fine-tuned on the curated dataset $\mathcal{D}_{\mathrm{adaptive}}$ using standard supervised fine-tuning (SFT). The learning objective is to maximize the likelihood of generating the optimal reasoning chain $\mathrm{CoT}_{i}^{\mathrm{opt}}$ and the correct answer $y_{i}$ given the input question $x_{i}$:
$\mathcal{L}_{\text{adaptive}}=-\sum_{(x,y,\text{CoT}^{\text{opt}})\in\mathcal{D}_{\text{adaptive}}}\sum_{t}\log P(w_{t}|x,w_{<t})$ (4)
where $w_{t}$ is the $t$-th token in the sequence $\mathrm{CoT}^{\mathrm{opt}},y$].
The Essence of Adaptive Learning. The key innovation is what the model learns from $\mathcal{D}_{\mathrm{adaptive}}$ : For simple problems that required only a short $\mathrm{CoT}^{\mathrm{opt}}$ in the dataset, the model learns to generate concise reasoning. For complex problems that required a longer $\mathrm{CoT}^{\mathrm{opt}}$, it learns to deploy more elaborate reasoning steps.
Unlike the model fusion step which controls behavior externally via the $\alpha$ parameter, $M_{\mathrm{adaptive}}$ learns to internalize the decision-making process. It does not merely imitate a fixed style but learns a spectrum of behaviors and, crucially, the conditional logic of when to apply each style. The training process teaches the model to approximate the function $f(x)=\mathrm{CoT}^{\mathrm{opt}}$, effectively compressing the entire model spectrum and selection pipeline into a single network.
4 Experiments
4.1 Experimental Setup
Datasets. We evaluate our method on five challenging mathematical reasoning benchmarks: GSM8K (Cobbe et al., 2021), MATH-500 (Lightman et al., 2023), AMC23 (Committees, 2023), OLYMPAID (He et al., 2024), and AIME25 (Committees, 2025). These datasets span a wide range of difficulty levels, providing comprehensive evaluation of reasoning capabilities.
Metrics. We employ three key metrics to comprehensively evaluate performance: (1) Pass@1: Problem-solving accuracy (primary quality metric); (2) ACT (Average chain-of-thought Tokens): Average number of tokens in the reasoning chain (efficiency metric); (3) AAT (Average Answer Tokens): Average total tokens in model output including reasoning and final answer (overall efficiency metric).
Baselines. We also present a comprehensive comparison between our proposed method and several state-of-the-art reasoning optimization approaches. We categorize the baseline approaches into the following groups: (1) Prompt-based methods: These rely on prompt engineering without parameter updates. We select CoD (Xu et al., 2025) and TALE-EP (Han et al., 2024) as our baselines. (2) RL-based methods: These leverage reinforcement learning, typically with reward signals derived from task-specific objectives or human feedback, to optimize reasoning performance. We select Z1 (Yu et al., 2025), DAST (Shen et al., 2025), AdaptThink (Zhang et al., 2025a) as our baselines. (3) SFT-based methods: These utilize supervised fine-tuning on curated datasets to learn reasoning patterns. We select AutoL2S (Luo et al., 2025) as our baseline.
Implementation Details. We obtained Long CoT data from DeepSeek-R1-Distill(tuanha1305, 2025) dataset, which contains problem-solution pairs with detailed reasoning chains, and the short CoT compressed from the long chains using DeepSeek-R1 model with specific compression prompts. The complete prompt is provided in Appendix LABEL:A:prompt. We use Qwen3-14B as our base model to train a distilled version for model fusion. The training details can be refered in Appendix LABEL:A:train_full and the model fusion details can be refered in Appendix A.1. Adaptive training data is constructed from the training splits of GSM8K (Cobbe et al., 2021) and MATH datasets using our proposed curation pipeline. Please refer to Appendix A.2 for adaptive training details.
4.2 Main Results
Table 1: Performance comparison of different methods across datasets.
Method GSM8K MATH-500 AMC23 OLYMPAID AIME25
Pass@1 ACT AAT Pass@1 ACT AAT Pass@1 ACT AAT Pass@1 ACT AAT Pass@1 ACT AAT
Qwen3-4B
Vanilla 95.2 1253.31 1557.70 96.0 5894.34 6699.73 97.5 10524.80 11362.50 72.9 12298.35 14863.27 70.0 16379.95 21496.90
Prompt-based
CoD 94.9 855.34 955.29 95.2 3676.47 4254.38 97.5 8535.40 9256.63 72.3 10065.13 12004.26 73.3 16226.83 20284.77
(-0.3) (-31.8%) (-38.7%) (-0.8) (-37.6%) (-36.5%) (+0.0) (-18.9%) (-18.5%) (-0.6) (-18.1%) (-19.2%) (+3.3) (-0.9%) (-5.6%)
TALE-EP 94.6 4294.88 5100.53 94.8 6533.66 8040.52 97.5 10773.84 12348.10 74.2 11947.04 14661.11 76.7 17718.96 21450.47
(-0.6) (+242.7%) (+227.4%) (-1.2) (+10.8%) (+20.0%) (+0.0) (+2.4%) (+8.7%) (+1.3) (-2.9%) (-1.4%) (+6.7) (+8.2%) (-0.2%)
SFT-based
DART(Ours) 93.9 401.13 596.37 96.4 3391.92 3981.97 100.0 6661.93 7379.20 72.0 8758.18 10271.33 80.0 16071.10 17513.30
(-1.3) (-68.0%) (-61.7%) (+0.4) (-42.5%) (-40.6%) (+2.5) (-36.7%) (-35.1%) (-0.9) (-28.8%) (-30.9%) (+10.0) (-1.9%) (-18.5%)
Qwen3-8B
Vanilla 95.7 1887.56 2214.61 94.4 4543.18 5309.38 92.5 8001.18 9436.85 68.6 9850.64 11257.47 56.7 15110.00 19063.27
Prompt-based
CoD 95.6 498.38 598.36 94.6 2881.84 3539.03 95.0 5949.68 6697.40 67.0 7713.47 9008.56 63.3 14399.34 15984.17
(-0.1) (-73.6%) (-73.0%) (+0.2) (-36.6%) (-33.3%) (+2.5) (-25.6%) (-29.0%) (-1.6) (-21.7%) (-20.0%) (+6.6) (-4.7%) (-16.2%)
TALE-EP 93.5 1148.60 1421.65 94.6 3617.01 4580.30 97.5 5987.89 7438.82 62.4 8637.41 10268.04 60.0 15473.76 19111.93
(-2.2) (-39.1%) (-35.8%) (+0.2) (-20.4%) (-13.7%) (+5.0) (-25.2%) (-21.2%) (-6.2) (-12.3%) (-8.8%) (+3.3) (+2.4%) (+0.3%)
SFT-based
DART(Ours) 95.1 983.87 1262.60 95.6 3321.36 3985.53 97.5 5204.95 5996.90 68.0 7468.58 8549.27 66.7 12610.43 13560.50
(-0.6) (-47.9%) (-43.0%) (+1.2) (-26.9%) (-24.9%) (+5.0) (-34.9%) (-36.5%) (-0.6) (-24.2%) (-24.1%) (+10.0) (-16.5%) (-28.9%)
Qwen3-14B
Vanilla 95.8 1399.16 1709.04 94.8 4075.58 4776.44 97.5 6691.5 7544.35 70.5 8695.07 10086.94 63.3 13324.23 16878.13
Prompt-based
CoD 95.9 535.22 617.18 95.4 2532.20 2988.71 97.5 4888.58 5563.3 70.2 6837.84 7685.47 66.7 12811.03 15066.03
(+0.1) (-61.7%) (-63.9%) (+0.6) (-37.9%) (-37.4%) (+0.0) (-26.9%) (-26.3%) (-0.3) (-21.4%) (-23.8%) (+3.4) (-3.9%) (-10.7%)
TALE-EP 92.8 94.10 169.78 79.2 368.34 655.64 70.0 1015.75 1424.80 48.0 764.35 2072.40 16.7 950.93 1617.63
(-3.0) (-93.3%) (-90.1%) (-15.6) (-91.0%) (-86.3%) (-27.5) (-84.8%) (-81.1%) (-22.5) (-91.2%) (-79.5%) (-46.6) (-92.9%) (-90.4%)
SFT-based
DART(Ours) 96.4 923.04 1165.95 96.4 3161.88 3748.81 100.0 4831.25 5601.65 70.4 7165.85 8206.90 70.0 11779.24 13446.47
(+0.6) (-34.0%) (-31.8%) (+1.6) (-22.4%) (-21.5%) (+2.5) (-27.8%) (-25.8%) (-0.1) (-17.6%) (-18.6%) (+6.7) (-11.6%) (-20.3%)
DeepSeek-R1-Distill-Qwen-7B
Vanilla 90.2 895.19 1007.26 91.0 2847.29 3385.94 90.0 5288.93 5789.63 57.8 6933.88 8003.48 36.7 13276.79 15060.97
Prompt-based
CoD 83.6 184.62 312.68 86.2 1587.01 1902.69 87.5 4383.23 4723.45 55.6 4792.10 5407.04 40.0 11009.93 11495.60
(-6.6) (-79.4%) (-69.0%) (-4.8) (-44.3%) (-43.8%) (-2.5) (-17.1%) (-18.4%) (-2.2) (-30.9%) (-32.4%) (+3.3) (-17.1%) (-23.7%)
TALE-EP 90.1 927.17 985.47 62.6 2809.62 2860.74 70.0 6459.59 6208.88 45.5 7583.87 8113.98 23.3 12607.29 12261.93
(-0.1) (+3.6%) (-2.2%) (-28.4) (-1.3%) (-15.5%) (-20.0) (+22.1%) (+7.2%) (-10.1) (+58.3%) (+50.1%) (-13.4) (-5.0%) (-18.6%)
RL-Based
Z1 89.3 - 591.38 74.6 - 1437.84 37.5 - 2657.4 37.8 - 2317.77 10.0 - 3986.63
(-0.9) - (-41.3%) (-16.4) - (-57.5%) (-52.5) - (-54.1%) (-20.0) - (-71.0%) (-26.7) - (-73.5%)
DAST 91.6 860.06 976.15 92.6 3123.93 3666.38 90.0 4657.95 5820.28 60.9 6860.06 7820.03 33.3 12727.72 13931.33
(+1.4) (-3.9%) (-3.1%) (+1.6) (+9.7%) (+8.3%) (+0.0) (-11.9%) (+0.5%) (+3.1) (-1.1%) (-2.3%) (-3.4) (-4.1%) (-7.5%)
AdaptThink 84.7 174.97 457.38 86.0 2226.04 2716.27 82.5 4882.35 5370.65 53.9 6538.94 7480.51 36.7 13891.56 16235.87
(-5.5) (-80.5%) (-54.6%) (-5.0) (-21.8%) (-19.8%) (-7.5) (-7.7%) (-7.2%) (-3.9) (-5.7%) (-6.5%) (+0.0) (+4.6%) (+7.8%)
SFT-Based
AutoL2S 91.0 - 481.34 77.6 - 1326.59 47.5 - 3639.50 41.9 - 3011.38 10.0 - 5303.93
(+0.8) - (-52.2%) (-13.4) - (-60.8%) (-42.5) - (-37.1%) (-15.9) - (-62.4%) (-26.7) - (-64.8%)
DART(Ours) 89.1 168.00 358.40 88.6 1853.43 2355.73 90.0 3460.26 4518.10 55.4 5076.86 6406.84 36.7 8729.72 9974.80
(-1.1) (-81.2%) (-64.4%) (-2.4) (-34.9%) (-30.4%) (+0.0) (-34.6%) (-22.0%) (-2.4) (-26.8%) (-19.9%) (+0.0) (-34.2%) (-33.8%)
Table 1 presents the comprehensive comparison across all datasets and model scales. Our proposed DART method demonstrates superior efficiency-accuracy trade-offs compared to all baselines.
Consistent Efficiency Gains.
DART consistently reduces computational cost across all model scales and datasets. The efficiency gains vary with problem difficulty: on simpler datasets like GSM8K, token usage is reduced by 34.0%–81.2%; while on highly challenging benchmarks AIME25, the method still achieves substantial savings up to 34.2%, demonstrating its ability to adaptively allocate more tokens to harder problems. This validates our core hypothesis that optimal reasoning length should adapt to problem difficulty.
Accuracy Preservation. DART maintains competitive accuracy across most settings and, notably, improves Pass@1 accuracy in several cases—such as on MATH-500, AMC23 and AIME25 (all Qwen3 scales)—while simultaneously reducing computational cost. This result suggests that adaptive termination can eliminate redundant or counterproductive reasoning steps.
4.3 Comparison to Prior Work
As shown in Table 1, our experimental analysis reveals a significant methodological divide in existing adaptive reasoning approaches. reinforcement learning (RL)-based methods (Z1, DAST, AutoL2S) demonstrate strong model specialization but are predominantly developed and evaluated exclusively on DeepSeek-R1-Distill-Qwen-7B, with no available implementations for Qwen3 series models. Conversely, prompt-based approaches (CoD, TALE-EP) offer broader model compatibility but face fundamental limitations in accuracy-efficiency trade-offs. This landscape highlights the unique positioning of our method in overcoming these constraints.
Superior Model Compatibility and Flexibility. Unlike RL-based methods that struggle with Qwen3’s extensive RLHF training, DART demonstrates remarkable architectural agnosticism. It achieves consistent improvements across all tested models (Qwen3-4B/8B/14B and DeepSeek-R1), proving its adaptability to diverse model architectures. While prompt-based methods like CoD and TALE-EP show wider model adaptability than RL-based techniques, they incur substantial accuracy costs across all model sizes. For instance, CoD exhibits accuracy degradation up to 6.6% on DeepSeek-R1, while TALE-EP shows similar patterns on Qwen3 series.
Enhanced Generalization Capabilities. DART exhibits superior cross-dataset generalization compared to both RL-based and prompt-based approaches. Despite being trained only on GSM8K and MATH, it maintains strong performance on out-of-distribution benchmarks (AMC23, OLYMPAID, AIME25). Prompt-based methods like CoD and TALE-EP show inconsistent generalization patterns, with significant accuracy variations across datasets. RL-based methods demonstrate even narrower generalization, often failing to transfer adaptive policies beyond their training distribution. DART’s supervised learning framework appears to capture more fundamental reasoning-length principles that translate better across problem types.
Training Stability and Practical Deployment. The supervised learning paradigm of DART offers significant advantages over RL-based methods in terms of training stability and reproducibility. While RL methods require careful reward engineering and suffer from optimization instability, DART employs standard fine-tuning procedures that yield consistent results. In contrast to methods needing intricate prompt design, DART’s trained approach ensures reliable inference performance without manual intervention. This makes it ideal for production environments that demand consistency.
The comprehensive comparison shows that DART achieves what previous methods cannot: effective adaptive reasoning across diverse model architectures with robust generalization and practical deployability.
4.4 Further Analysis
4.4.1 Validation of Model Fusion Effectiveness
Figure 3: Effect of fusion coefficient ($\alpha$) on compression ratio and accuracy.
In practice, we sample $11$ values of $\alpha$ (e.g., $\alpha\in\{0,0.1,...,0.9,1\}$) to create a discrete but dense ensemble of models $\{M_{\alpha_{1}},M_{\alpha_{2}},...,M_{\alpha_{11}}\}$. This ensemble allows us to approximate the continuum and explore the full range of reasoning lengths.
As shown in Fig. 3, we observe that the average length of the generated CoT is a smooth, monotonically decreasing function of $\alpha$. This confirms that our fusion method successfully creates a controllable knob for adjusting reasoning complexity without additional training. Furthermore, the accuracy on reasoning benchmarks reveals a more nuanced relationship: as $\alpha$ decreases from 1 to 0, accuracy initially shows a slight improvement before entering a declining phase. This non-monotonic behavior demonstrates that longer reasoning chains do not invariably lead to higher accuracy; beyond a certain point, excessive verbosity can be counterproductive. The fusion spectrum thus effectively samples different points on the efficiency-accuracy Pareto frontier, providing an efficient and elegant solution for generating the multi-length reasoning data required for the next step of our pipeline.
4.4.2 Impact of Data Curation Repetition
The adaptive training data generation process involves multiple passes through our model spectrum to identify optimal reasoning chains. Fig. 4 presents the effect of repetition frequency on final model performance:
Consistency Across Repetitions. Performance remains stable across passes. For Qwen3-4B on MATH-500, accuracy stays high (95.4%–96.4%) with only minor ACT variation (3355.4–3696.0). Similarly, for Qwen3-8B on GSM8K maintains  95% accuracy across passes. This indicates that a single model pass (Pass@1) is sufficient to collect high-quality adaptive examples.
Efficiency-Accuracy Trade-offs. Additional passes bring limited and inconsistent efficiency gains. For example, Qwen3-4B on GSM8K shows ACT reduction from 512.93 to 401.13 over three passes, but this trend does not hold on harder datasets. This suggests diminishing returns beyond the first pass, making Pass@1 the most practical choice.
Figure 4: Effect of data curation iterations on model performance across different datasets.
4.4.3 Sensitivity to $\alpha$ Sampling Density.
We compare loose (5 points), middle (10 points), and dense (20 points) $\alpha$ sampling strategies in Table 2:
Density-Accuracy Relationship. Contrary to expectations, denser sampling does not consistently improve accuracy. On Qwen3-8B/MATH-500, loose sampling achieves the highest accuracy (96.6%), while dense sampling gives comparable results (96.0%). A sparse sampling (5–10 points) thus sufficiently captures key reasoning behaviors.
Efficiency Optimization. Efficiency, however, improves with density: on Qwen3-4B/GSM8K, dense sampling yields the lowest ACT (421.67) versus loose (986.11) and middle (725.80), though with a slight accuracy drop (93.6% vs. 95.5%). Loose sampling favors accuracy; dense sampling favors token reduction.
Based on these findings, we recommend middle-density sampling (10 $\alpha$ points) as the default configuration, balancing computational cost during data curation with final model performance. This density provides sufficient granularity to identify near-optimal reasoning lengths without excessive computational overhead.
Table 2: Impact of $\alpha$ sampling density on model performance across different datasets.
Sampling GSM8K MATH-500 AMC23 OLYMPAID AIME25
Density Pass@1 ACT AAT Pass@1 ACT AAT Pass@1 ACT AAT Pass@1 ACT AAT Pass@1 ACT AAT
Qwen3-4B
Loose 95.5 986.11 1230.04 95.6 3613.50 4248.40 100.0 6224.53 6929.80 71.4 9322.56 11327.69 76.7 14323.03 16386.83
Middle 94.7 725.80 952.60 95.6 3344.17 3893.69 100.0 6824.88 7590.55 70.8 9151.90 10623.66 70.0 14345.74 17057.70
Dense 93.6 421.67 602.24 96.2 3637.91 4166.00 100.0 5803.83 6409.65 71.1 9392.36 10385.10 76.7 15556.56 18016.97
Qwen3-8B
Loose 95.7 1114.57 1397.32 96.6 3377.11 3974.48 97.5 5503.95 6282.48 67.6 7881.85 8952.07 70.0 12986.69 14652.53
Middle 95.7 1095.47 1408.25 95.4 3271.15 3911.19 100.0 6412.08 7243.20 67.9 7744.78 8797.41 63.3 12235.40 15225.40
Dense 94.8 682.98 924.61 96.0 3572.20 4172.57 95.0 5719.13 6323.32 68.6 7722.53 8741.07 70.0 12546.3 13584.63
5 Conclusion
In this work, we presented DART, a supervised framework for difficulty-adaptive reasoning truncation that allows LLMs to dynamically adjust their thinking length according to problem complexity. By distilling concise reasoning patterns, interpolating them into a continuum of reasoning styles, and curating optimal training signals, it learns when to stop thinking. DART realizes significant reasoning truncation and speedup without sacrificing accuracy, paving the way toward more efficient and sustainable LLMs.
Appendix A Appendix
A.1 Details of Model Fusion
We use Qwen3-14B as our base model to train a distilled version for model fusion. The fusion coefficient ($\alpha$) we used can be referred in Table 3.
Table 3: Fusion coefficient ($\alpha$) used for model fusion.
Dataset Alpha Accuracy Avg Tokens Ratio
ACT AAT
GSM8K 0 (Qwen3-14B) 96.6 1325.15 1632.77 100.00%
0.05 96.5 1288.49 1589.59 97.23%
0.1 96.4 1206.89 1496.72 91.08%
0.125 96.6 1141.44 1425.52 86.14%
0.15 96.5 1108.54 1400.72 83.65%
0.175 96.4 1002.43 1263.96 75.65%
0.2 96.4 988.16 1252.73 74.57%
0.225 96.7 957.73 1199.52 72.27%
0.24 96.6 951.57 1193.46 71.81%
0.25 96.5 846.61 1042.58 63.89%
0.275 96.5 741.08 918.67 55.92%
0.3 96.6 703.36 882.84 53.08%
0.325 96.4 668.81 844.13 50.47%
0.35 96.5 647.15 825.05 48.84%
0.375 96.5 575.50 743.11 43.43%
0.4 96.6 487.18 657.17 36.76%
0.42 96.2 409.42 578.93 30.90%
0.44 96.2 360.85 530.66 27.23%
0.46 96.2 308.48 477.95 23.28%
0.48 96.0 279.91 444.97 21.12%
0.49 96.1 278.02 447.05 20.98%
0.495 96.2 275.14 444.18 20.76%
0.498 96.1 272.52 437.76 20.57%
0.5 95.9 110.27 284.48 8.32%
0.6 95.6 105.16 285.95 7.94%
0.7 95.4 104.50 287.29 7.89%
0.8 95.4 102.56 291.24 7.74%
0.9 94.9 102.59 288.54 7.74%
1.0 (Distilled) 94.7 102.45 297.50 7.73%
MATH 0 (Qwen3-14B) 95.4 4109.18 4865.18 100.00%
0.1 95.4 3924.72 4629.28 95.51%
0.2 95.6 3572.30 4229.31 86.93%
0.25 95.2 3347.99 3895.70 81.48%
0.275 94.9 3089.26 3593.51 75.18%
0.3 95.0 2978.05 3488.89 72.47%
0.325 94.5 2813.51 3299.48 68.47%
0.35 94.6 2756.00 3246.53 67.07%
0.375 93.9 2500.21 2972.72 60.84%
0.4 92.6 2227.13 2697.19 54.20%
0.42 91.9 2047.34 2508.61 49.82%
0.44 91.3 1911.97 2373.67 46.53%
0.46 90.5 1763.34 2223.73 42.91%
0.48 90.3 1697.39 2168.32 41.31%
0.49 90.0 1675.61 2163.59 40.78%
0.495 90.2 1677.61 2121.31 40.83%
0.498 89.1 1660.45 2125.71 40.41%
0.5 79.8 421.82 888.41 10.27%
0.6 77.1 303.95 775.21 7.40%
0.7 76.3 284.95 736.22 6.93%
0.8 73.4 261.47 733.30 6.36%
0.9 71.6 247.82 713.24 6.03%
1.0 (Distilled) 70.5 244.44 686.38 5.95%
A.2 Details of Adaptive Model Training
We perform adaptive fine-tuning on the model using the LLaMA-Factory (Zheng et al., 2024) framework with Low-Rank Adaptation (LoRA). The LoRA configuration targets all linear layers (lora_target: all) with a rank of 256 and an alpha of 16. The training source question data comprises about 15,000 samples from the GSM8K and MATH mixture dataset, using the qwen3 prompt template. A cutoff length of 32,768 tokens is applied to all sequences.
The model is trained for 3 epochs using the adamw_torch optimizer and the following hyperparameters:
  • Cutoff length: 32,768
  • Max samples: 15,000
  • Batch size: 1 (with gradient accumulation of 8)
  • Learning rate: $2\times 10^{-5}$ with a cosine schedule and a warmup ratio of 0.1
  • Precision: bf16
  • LoRA rank: 256
  • LoRA alpha: 16
  • Validation split: 10% of the training data
  • Evaluation strategy: every 200 steps
The training is accelerated with performance optimizations, including the Liger kernel and Unsloth’s garbage collector. All experiments utilize 16 parallel workers for data preprocessing and are configured with overwrite_cache=true.
A.3 Details of Evaluation
We use the Qwen2.5-Math (Yang et al., 2024) framework for unified evaluation across tasks. In our evaluation setup, all models were constrained to a maximum generation length of 32,768 tokens and temperature of 0.6 to align with DeepSeek’ technical report (Guo et al., 2025) and Qwen3’ technical report (Yang et al., 2025).